OpenStudy (anonymous):

Which of the following is the solution to the equation 25^(z – 2) = 125 ?

jimthompson5910 (jim_thompson5910):

Hint: 25 = 5^2 and 125 = 5^3 ------------------- So 25^(z – 2) = 125 (5^2)^(z – 2) = 5^3 5^(2(z-2)) = 5^3 Now equate the exponents to get 2(z-2) = 3

jimthompson5910 (jim_thompson5910):

I'll let you finish up and solve for z

OpenStudy (anonymous):

7/2?

jimthompson5910 (jim_thompson5910):

that is 100% correct

OpenStudy (anonymous):

z = 3.5

jimthompson5910 (jim_thompson5910):

yes, 7/2 = 3.5, so they are the same thing